# Optimization of deterministic photonic graph state generation via local operations

### v1

@@ Abstract
Realizing photonic graph states, crucial in various quantum protocols, is challenging due to the absence of deterministic entangling gates in linear optics. To address this, emitter qubits are leveraged to establish and transfer the entanglement to photons. We introduce an optimization method for such protocols based on the local Clifford equivalency of states and the graph-shape correlated generation cost parameters. Employing this method, we achieve a 50% reduction in use of the 2-qubit gates for generation of the repeater graph states and a 65% reduction in the total gate count for 15-node random dense graphs.
@@
Figure 1: Schematic circuit for sequential generation of photonic graph states. The three main components are (i) single qubit gates on photons (yellow), (ii) emitter blocks representing arbitrary operations on the set of emitter qubits between emission events (green) , and (iii) emission events (blue) corresponding to a two-qubit emitter-photon gate.
@@
Devising the necessary operations between emission events for generating a target state is not trivial. As qubit connectivity is heavily constrained, this task is fundamentally different from a general quantum circuit synthesis problem of establishing a desired quantum channel [21, 22]. To date, only a limited number of resource-conscious generation protocols have been proposed, with the majority designed for specific categories of graphs [16, 23, 24, 25] and only a few for arbitrary states [26, 27], with the optimization efforts primarily concerned with minimizing the number of emitters. Although the emitter count is a significant factor, there are other important metrics that can be considered in the circuit design, such as the required emitter coherence times and the number of two-qubit gates.
In this paper, we propose a method to reduce the generation costs of photonic graph states, by leveraging the concept of local Clifford (LC) equivalency of graphs [28]. While LC equivalency has been previously employed to optimize the preparation of graph states [22, 29, 30], this is the first time that it is applied to the case of deterministic photonic-state generation, differing notably from non-photonic cases. The method works through identifying a set of alternative intermediate graphs which can be converted to the final target state with minimal overhead, providing a selection of generation routes to choose from. Moreover, we adopt a novel approach by employing graph theoretical metrics to correlate the structural characteristics of a graph with the cost parameters of its generating circuit, such as circuit size and depth. This is used to determine which intermediate graphs are more favorable for the process, filtered solely based on their shape. The use of LC equivalency, combined with the correlations between graph shape and circuit resources, forms the basis of a versatile optimization framework for photonic state generation.
We demonstrate that a correlation between the circuit size and the edge count within the local equivalency classes of dense random graphs leads to average reductions of 65% and 46% in circuit gate count and ‘emitter depth’ (to be defined below), respectively, for 15-node graphs. In another case, an analysis of the repeater graph states (RGS), crucial resources for quantum communication protocols [9], resulted in the identification of the optimized elementary circuit blocks for the generation of an arbitrarily large RGS with a 50% reduction in the number of required two-qubit gates between emitters. Moreover, a polynomial time algorithm to generate the full local equivalency class of RGS is presented for the first time.
@@
Notable circuit parameters to consider in a cost analysis are (i) the number of emitters, (ii) the number of gates of each type, and (iii) the circuit depth. In particular, using more emitters necessitates coherent control over a larger set of qubits, and a wider extent of emitter-emitter couplings, which is challenging in practice. This emitter coupling cost can be directly reflected in the number of two-qubit gates which, apart from being difficult to implement, are also a primary source of infidelity in the circuit. The depth of the circuit is associated with the maximum number of successive gates applied on the qubits and is also indicative of the time it takes to complete the protocol; two significant factors in circuit performance. Since an emitter can undergo a reset after each measurement, as seen in protocols outlined in Refs. [27, 24], and its coherence time can be considered to reset as well, a more physically relevant metric would be the maximum number of consecutive gates on an emitter in the interval between each reset to |$\ket{0}$ and a subsequent measurement. Here we define this new metric as ‘emitter depth’, complementary to the regular circuit depth.
In our analysis, apart from the emitter count, the focus is on other experimentally significant metrics that have been mostly overlooked in the existing literature on deterministic photonic state generation, e.g., the number of emitter-emitter cnots, and the emitter depth. Nevertheless, the optimization process can be easily adapted to incorporate other customized cost functions.
@@
The LC orbit can be leveraged for cost reduction whenever a deterministic graph-to-circuit map, such as the protocol in Ref. [27], is employed to find a generation circuit for a given graph. Since any two graphs within the same orbit can be converted to one another by local unitaries, i.e., single-photon gates, and given that–as previously argued–such gates have no significant contribution to our circuit cost, the minimum generation cost for all orbit members is identical and is determined by the orbit member that yields the best circuit under the utilized graph-to-circuit mapping protocol. The goal is thus to identify and generate a more favorable alternative graph in the orbit and then convert back to the original target.
@@
Another set of alternative targets can also be obtained by permuting the temporal emission order of the photons which provides a natural way of labelling the nodes in a graph. All isomorphic graphs found by changing the photon emission order share the same entanglement structure and thus can be considered identical once generated, conditioned by photon indistinguishability. The generation cost, however, varies from case to case as generating circuits depend on emission order [27].
@@
To compile an exhaustive array of alternative target graphs, the LC orbit is first identified via iterative application of the LC rule to graph nodes until no new graphs can be found. Subsequently, all possible emission orderings are considered for each orbit member. The set of alternative targets is then fed to a graph-to-circuit map in order to create a list of generating circuits, which are then analyzed to find the best option(s) with respect to a chosen performance metric. For our choice of graph-to-circuit map, we make use of the protocol outlined in Ref. [27]. This mapping protocol guarantees the use of a minimal number of emitters when emission order is fixed, and runs in polynomial time, making it a suitable benchmark for evaluating cost reduction performance of our proposed method. Hereafter, comparisons are made between the optimized and original cases with the ‘original’ circuit obtained using this protocol for a given graph.
@@
Results.—In general, finding the exhaustive set of alternative graphs is not always feasible except for small-sized or highly symmetric cases. This is because the LC orbit can grow rapidly with the size of the graph; for instance, for linear and cycle graphs the scaling is exponential [33]. In fact, determining the size of the orbit of an arbitrary graph is a #P-complete problem [36]. Due to this complexity, previous investigations into the LC orbit have been limited to graphs with a maximum of 12 nodes [22, 35]. Furthermore, the number of isomorphic graphs found by permutation of the emission order scales factorially with size, making the exhaustive search impractical for large graphs.
Here, an exhaustive analysis is conducted, considering both local equivalency and isomorphism classes of graphs with up to 7 qubits [37], enumerating the 45 entanglement classes outlined in Ref. [32]. For each class the most efficient generating circuits are identified, providing recipes for the generation of any of the $\sim 1.9\times 10^{6}$ labelled connected graphs falling under this category [38]. All the analyses presented here and below were conducted using our Python package “GraphiQ” [39], developed for simulating, manipulating, and evaluating photonic graph states and their generating circuits, run on a 3.22 GHz, 8-core, Apple M1 Pro chip with 16 GB of RAM.
While a full orbit analysis is generally impractical, sampling it is feasible by applying LC operations on a random sequence of nodes. This sample can then be used to identify correlations between graphs and their generating circuits, which can serve as guidelines for an oriented search within the orbit to reach lower-cost graphs. We remark that the type and degree of the correlations are specific to each individual orbit and also depend on the selected graph-to-circuit mapping protocol.
@@
As a demonstration, the correlation between the graph edge count and its circuit size was investigated for an orbit sample of size 100 for random Erdös-Rényi graphs [40] with $n$=15 nodes and $0.95$ edge probability. Averaging over 1000 instances, the Pearson correlation coefficient [41]–a number between $-1$ to $1$ with $1$ indicating a positive linear correlation–was found to be $0.83\pm 0.10$ for the number of unitaries and $0.73\pm 0.14$ for the emitter depth, both serving as circuit size indicators. Such correlations can then be employed to reduce the generation cost; specifically, we utilized an efficient edge reduction strategy (see Supplemental Material [42]) via LC operations to actively aim for the graphs in the orbit that are more likely to have a lower generation cost. For graphs with $n$=15, reductions of $65\%\pm 11$% in the number of unitaries and $46\%\pm 18$% in emitter depth were obtained by averaging over 1000 cases. Fig. 3 shows the cost reduction achieved when extending the use of this strategy to larger sizes.
Figure 4: (Top) Original RGS vs the best alternative shape in its orbit. (Bottom) The elementary circuit blocks associated with the above graphs; the implementation of the block adds two photons to the state under preparation. The alternative block features a reduced number of cnots. $U$ and $U^{\prime}$ are Clifford unitaries $HPHPX$ and $HPHP$, respectively.
Compared to a random search, the correlation-based approach does not require analysis of all the graphs and their circuits in a sample. Instead, it leverages an oriented navigation of the orbit, enabling a substantial reduction in optimization runtime. Specifically, when comparing a random search method to the use of an edge reduction strategy for graph sizes in the range of $n$=30 to 60, we observed a one to two orders of magnitude reduction in runtime while maintaining a similar performance. Refer to the Supplemental Material [42] for a more thorough discussion of this procedure.
We remark that for structured graphs, i.e, graphs with shapes that can be parameterized by size, it is possible to perform a full orbit analysis for smaller graphs to gain insight and identify general patterns that may be applicable to the larger-sized versions of them. One example of interest for such cases is the RGS. An RGS performs as an all-photonic quantum repeater and is of significant importance in facilitating the realization of long distance quantum communications [24, 43, 44]. By exploring the orbit of small cases of RGS, through a dynamical system mapping, we achieved a 50% reduction in the number of emitter-emitter cnots and were able to generalize this result for the RGS with arbitrary $n$ qubits (see Fig. 4), i.e., reducing the cnot count from $n-3$ to $n/2-2$. Specifically, as shown in the correlation diagram of Fig. 5, we define the variables of the dynamical system as the edge-count of a graph (x-axis) and the cnot count of its corresponding circuit (y-axis) for all graphs in the orbit. By tracking the changes within the correlation diagram when LC transformations were applied, we were able to not only fully navigate the known orbit of a small RGS, but also to generalize the pattern and develop, for the first time, an efficient, polynomial-time algorithm to generate the orbit of an arbitrarily large RGS; the algorithm is available in the Supplemental Material [42]. It is noteworthy that the LC orbit identification task via brute-force search methods is not an efficient and scalable process in general. With a full knowledge of the orbit, we can identify the generalized graph shape associated with the least generation cost (e.g., here the number of cnots) and determine its elementary circuit block, to be employed to lower the overall generation cost of an RGS of any size. As evident in Fig. 4, the optimal block uses one cnot gate to add two photons to the state whereas in the original case, two cnots are required, enabling the change of scaling from $n$ to $n/2$. This optimized scaling of two-qubit gates, identified systematically here, aligns with the scaling of the ad-hoc protocol described in Ref. [24], which was specifically designed for the optimal generation of RGS using a different approach and a different emission sequence. Additional information on the RGS orbit can be found in the Supplemental Material [42].
Figure 5: Number of cnots vs Number of edges correlation diagram of 100-qubit RGS. Points correspond to graphs in the orbit, with the red point being the original RGS. The inset shows the portion of the orbit with low edge counts.
@@
Conclusion.—We introduce an optimization framework tailored for the deterministic generation of photonic graph states, a task that poses significant theoretical and experimental challenges due to its inherent constraints. Our approach employs local complementation to effectively identify alternative graphs that can serve as more resource-efficient avenues for reaching a target graph.
@@
In our correlation analyses, we establish that specific graph metrics can be linked to some cost parameters, enabling us to manage the use of resources at an orbit search level; a 50% reduction in the requisite number of emitter-emitter gates showcases one of the potential applications of this method. Yet, the door is open for exploring other graphs of interest with more elaborate graph theoretical metrics.
@@
It is worth noting that this optimization approach is independent of the graph-to-circuit mapping protocol employed, adding a flexible layer of applicability across existing state generation protocols. Moreover, the employed cost metrics can be tailored to meet the requirements of specific experimental platforms, e.g., as we have done with the ‘emitter depth’ to address the limitations in emitters’ coherence time. Additionally, the application can be easily extended to the probabilistic fusion-based generation protocols, as in Ref. [5], by facilitating the expansion of both the size scale and diversity of resource states used as building blocks.
In summary, our optimization framework presents a versatile method to identify resource-efficient graph state preparation schemes, paving the way for future advancements in quantum communication and computing technologies.

### v2

@@ Abstract
Realizing photonic graph states, crucial in various quantum protocols, is challenging due to the absence of deterministic entangling gates in linear optics. To address this, emitter qubits have been leveraged to establish and transfer the entanglement to photons. We introduce an optimization method for such protocols based on the local Clifford equivalency of states and the graph theoretical correlations of the generation cost parameters. Employing this method, we achieve a 50% reduction in use of the 2-qubit gates for generation of the arbitrary large repeater graph states and similar significant reductions in the total gate count for generation of random dense graphs.
@@
Figure 1: Schematic of the circuit used for the sequential generation of photonic graph states. The main components are (i) single qubit gates on photons (yellow), (ii) emitter blocks representing arbitrary operations on the set of emitter qubits between emission events (green) , and (iii) emission events (blue) corresponding to a two-qubit emitter-photon gate.
@@
Devising the necessary operations between emission events for generating a target state is not trivial. As qubit connectivity is heavily constrained, this task is fundamentally different from a general quantum circuit synthesis problem of establishing a desired quantum channel [21, 22]. To date, only a limited number of resource-conscious generation protocols have been proposed, with the majority designed for specific categories of graphs [16, 23, 24, 25] and only a few for arbitrary states [26, 27, 28], with the optimization efforts primarily concerned with minimizing the number of emitters. Although the emitter count is a significant factor, there are other important metrics that can be considered in the circuit design, such as the required emitter coherence times and the number of two-qubit gates.
In this paper, we propose a method to reduce the generation costs of photonic graph states, by leveraging the concept of local Clifford (LC) equivalency of graphs [29]. While LC equivalency has been previously employed to optimize the preparation of graph states [22, 30, 31], this is the first time that it is applied to the case of deterministic photonic-state generation, differing notably from non-photonic cases. The method works through identifying a set of alternative intermediate graphs which can be converted to the final target state with minimal overhead with respect to the photonic platform constraints, providing a selection of generation routes to choose from. While deterministic generation protocols such as Ref. [27] use Clifford circuits to generate a single fixed graph (the target), our approach differs by aiming to identify the best state to be generated, among all graphs in the local Clifford equivalency class. Any deterministic graph-to-circuit map can later be used to obtain the generation circuit. Moreover, we adopt a novel approach by employing graph theoretical metrics to correlate the structural characteristics of a graph with the cost parameters of its generating circuit, such as circuit size and depth. This is used to determine which intermediate graphs are more favorable for the process, filtered solely based on their shape. The key insight of this work is leveraging the size and correlations of the LC equivalency class, leading to efficient and substantial reductions in generation costs. This forms the basis of a versatile optimization framework for photonic state generation.
A primary result is the identification of the optimized elementary circuit blocks for generating arbitrarily large repeater graph states (RGS) [9], achieving a 50% reduction in required two-qubit gates between emitters. This reduction promises a potential orders-of-magnitude increase in the success rate of quantum repeater protocols. Moreover, we present for the first time a polynomial-time algorithm to generate the full local equivalency class of an RGS. We elaborate on the wide-range applicability of the new method by showcasing an efficient correlation-oriented optimization within the local equivalency classes of random dense graphs. We achieve average reductions of 65% and 46% in unitary gate count and ‘emitter depth’ (defined below), respectively, in generation circuits of 15-node graphs, and observe similar advantages as the graph size increases.
@@
Notable circuit parameters to consider in a cost analysis are (i) the number of emitters, (ii) the number of gates of each type, and (iii) the circuit depth. In particular, using more emitters necessitates coherent control over a larger set of qubits, and a wider extent of emitter-emitter couplings, which is challenging in practice. This emitter coupling cost can be directly reflected in the number of two-qubit gates which, apart from being difficult to implement, are also a primary source of infidelity in the circuit. The depth of the circuit is associated with the maximum number of successive gates applied on the qubits and is also indicative of the time it takes to complete the protocol; two significant factors in circuit performance. Since an emitter can undergo a reset after each measurement, as seen in protocols outlined in Refs. [27, 24], and its coherence time can be considered to reset as well, a more physically relevant metric would be the maximum number of consecutive gates on an emitter in the interval between each reset to |$\ket{0}$ and a subsequent measurement. Here, we define this new metric as ‘emitter depth’, complementary to the regular circuit depth.
In our analysis, while maintaining the minimum emitter count, the focus is on other experimentally significant metrics that have been mostly overlooked in the existing literature on deterministic photonic state generation, e.g., the number of emitter-emitter cnots, and the emitter depth. Nevertheless, the optimization process can be easily adapted to incorporate other customized cost functions.
@@
In this paper, we will show that the LC orbit can be leveraged for cost reduction whenever a deterministic graph-to-circuit map, such as the protocol in Ref. [27], is employed to find a generation circuit for a given graph. Since any two graphs within the same orbit can be converted to one another by local unitaries, i.e., single-photon gates, and given that–as previously argued–such gates do not significantly contribute to our circuit cost, the minimum generation cost for all orbit members is identical and is determined by the orbit member that yields the best circuit under the utilized graph-to-circuit mapping protocol. The goal is therefore to identify and generate a more favorable alternative graph in the orbit and then convert back to the original target.
@@
Another set of alternative targets can also be obtained by permuting the temporal emission order of the photons, which provides a natural way of labelling the nodes in a graph. All isomorphic graphs found by changing the photon emission order share the same entanglement structure and thus can be considered identical once generated, conditioned by photon indistinguishability. The generation cost, however, varies from case to case as generating circuits depend on emission order [27].
@@
To compile an exhaustive array of alternative target graphs, the LC orbit is first identified via iterative application of the LC rule to graph nodes until no new graphs can be found. Subsequently, all possible emission orderings are considered for each orbit member. The set of alternative targets is then fed to a graph-to-circuit map in order to create a set of generating circuits, which are then analyzed to be ranked with respect to a chosen figure of merit. Our method takes the graph-to-circuit map as an input but operates at a higher level to expand the set of targets. While our proposed optimization scheme is mapping protocol-agnostic, for concreteness, we have first focused on the protocol in Ref. [27] in our analyses. This protocol guarantees the use of a minimal number of emitters for a given photon emission order, and runs in polynomial time, making it a suitable benchmark for evaluating cost reduction performance. Hereafter, for a given graph, comparisons are made between the original circuit obtained from [27]’s protocol and its optimized counterparts. Lastly, to showcase that the framework can accommodate any graph-to-circuit mapping, we also implement the protocol from Ref. [28] in the optimization, obtaining similar advantages to those achieved with Ref. [27].
Results.
Figure 3: (a): Original RGS. (b, c): Partially transformed graphs corresponding to diagram (d)’s correlated part. (d): Number of cnots vs Number of edges for graphs in the orbit of a 100-qubit RGS. The red point is the original RGS. A linear trend is visible within a portion of the orbit. (e, f): The elementary circuit blocks to be applied on the 2 emitters required to generate graphs seen in (b) or (c), adding the photon pair $p_{2k}$ and $p_{2k+1}$ to the state under generation. Emissions are modeled by the cnots between an emitter and a photon. From (e) to (f), the rate of use of emitter-emitter cnots (shaded in circuit) changes from $1$ per photon to $1/2$ per photon emitted. (e) Generates a photon pair as in the repeated part of the original RGS structure, that is, a root and a leaf. (e) Generates a photon pair corresponding to a pair in the transformed part of the graph, indicated by the shaded area in (c). $H$, $P$, and $X$ gates are Hadamard, Phase, and Pauli $X$, and $U$ is the Clifford unitary $HPHPX$. The last operations in both blocks are a classically controlled measurement followed by a reset.
@@ Results.
In general, finding the exhaustive set of alternative graphs is not always feasible except for small-sized or highly symmetric cases. This is because the LC orbit can grow rapidly with the size of the graph. For instance, for linear and cycle graphs the scaling is exponential [34]. In fact, determining the size of the orbit of an arbitrary graph is a #P-complete problem [37]. Due to this complexity, previous investigations into the LC orbit have been limited to graphs with a maximum of 12 nodes [22, 36].
We remark that for structured graphs, which can be parameterized by size, a full orbit analysis of smaller graphs can lead to identifying general patterns that are applicable to the larger graphs of the same type. One example of interest is the RGS, shown in Fig. 3(a), which enables all-photonic quantum repeater protocols and facilitates long-distance quantum communications [24, 38, 39]. Exploring the orbit of small RGSs shows that a near 50% reduction in the number of emitter-emitter cnots is obtainable (see Fig. 3). In Fig. 3(d), we plot the correlation diagram for the circuit cnot count as a function of the edge count of their respective graphs in the orbit for a 100-qubit RGS. The orbit can be divided into a linearly correlated region and an uncorrelated one (shown in the inset) with respect to the edge count. Figure 3 (b) and (c) correspond to the first two graphs belonging to the linearly correlated part. To obtain the graphs shown in (b) and (c), local complementations are applied on the nodes of the original RGS shown in (a) in order of $\{1,3,1\}$ and $\{1,3,5,7,1\}$, respectively. Here, the node labeling (starting from 00) also corresponds to the photon emission order. Similarly, to obtain the $j$-$th$ graph along the linearly correlated part ($j=0$ being the original RGS) we must apply the LC rule on the odd numbered nodes up to $14j-$ 1, followed by an additional LC on node $1$. Remarkably, breaking such an LC rule sequence at any node, results in a unique graph. By comparing the sequences leading to two consecutive graphs ($j$ and $j+1$) we observe that the $j+1$-$th$ graph has two extra nodes, $14j+$ 1 and $34j+$ 3, before the final LC on node 1 in its sequence compared to the $j$-$th$ one. Breaking the sequence at these two points correspond to two graphs belonging to the uncorrelated region of the orbit, constituting the inset of Fig. 3(d). Such generalization allows us to navigate the orbit to reach the graph with the minimum cnot count. This also provides, for the first time, an efficient polynomial-time algorithm to generate and access the entire orbit of an arbitrarily large RGS. (See supplementary section (A) [40].) It is noteworthy that the LC orbit identification task via brute-force search methods is not an efficient and scalable process in general. The generation circuit blocks in Fig. 3(d-e) demonstrate how such graph transformations can change the scaling of the required number of 2-qubit gates from $n$ to $n/2$ for an $n$-qubit RGS, as the emission of each photon pair belonging to the transformed part requires a single cnot instead of the original two. The final optimal graph shape can be achieved by sequentially applying the LC rule to all central nodes of an RGS for odd $n/2$ (i.e., $\{1,3,\ldots,n-1\}$). For even $n/2$, the last node in the sequence is replaced by its adjacent leaf, resulting in the set $\{1,3,\ldots,n-2\}$. The cnot count is as a result reduced from $n-3$ to $n/2-2$. This optimized scaling, identified systematically here, aligns with the scaling of the ad-hoc protocol described in Ref. [24], which was specifically designed for the optimal generation of RGS using a different approach and a different emission sequence. Furthermore, employing the mapping protocol of Ref. [28] instead of Ref. [27], results in similar scaling advantages in terms of circuit depth. (See Supplementary section (B) [40].)
The importance of our results lies in the ability to double the maximum size of the producible RGS, which is limited by the emitters’ finite coherence time, as the cnots between emitters are the primary time-consuming operations in such circuits [28]. The significance of this size increase is better reflected in the repeater protocol’s success probability in establishing an EPR pair between two distant parties, which is given by $P=(1-(1-P_{b})^{n/4})^{N}$ [9]. Here, $n$ is number of photons in RGS, $N$ is the number of RGS stations between the two parties, and $P_{b}$ is the probability of a successful entanglement swapping between two photons from two adjacent RGSs. Doubling the size from $n$ to $2n$, especially in the realistic long distance regime with $n$ being in the order of tens and $N$ in hundreds, results in several orders-of-magnitude increase in the overall success rate. (See Supplementary section (C) [40] for an example.)
@@
To demonstrate the applicability of our method beyond structured graphs, we also investigated a set of graphs on the opposite end of the spectrum, i.e., random graphs [41]. Here, we limit the study to dense graphs with a 95% edge probability. Conjecturing a correlation between the generation cost and the shape of a graph, we chose a simple metric to parameterize the graph shape: the number of edges. This correlation can be directly confirmed for graphs of small size, e.g., 15 nodes, as the whole orbit is accessible. By choosing the graph with the least edge count in the orbit, we observed average reductions of $65\%\pm 11\%$ in the number of unitaries and $46\%\pm 18\%$ in emitter depth. One can efficiently test this correlation for larger graphs, as there are polynomial-time methods to find sparse graphs within the orbit of dense graphs (see supplementary section (D) [40] for one such strategy developed by the authors). Figure 4 shows the cost reduction achieved when extending the use of this strategy for graph of up to 65 nodes.
We also conducted an exhaustive analysis (see Supplementary Section E [40]) of graphs with up to 7 qubits [42], enumerating all the possible 45 entanglement classes as defined in Ref. [33] and comparing the generation circuits across $\sim 1.9\times 10^{6}$ labeled connected graphs [43]. All the presented analyses were conducted on a 3.2 GHz, 8-core, Apple M1 Pro chip with 16 GB of RAM, using our Python package “GraphiQ” [44].
@@
Discussion.—We introduce an optimization framework tailored for the deterministic generation of photonic graph states, a task that poses significant theoretical and experimental challenges due to its inherent constraints. Our approach leverages the concept of local Clifford equivalency and local complementation to effectively identify alternative graphs in the same local equivalency class, that can serve as more resource-efficient paths for reaching a target state.
@@
In our correlation analyses, we establish that specific graph metrics can be linked to circuit cost parameters, enabling us to manage the use of resources at an orbit search level. Future work should explore other graph structures of interest and with more elaborate graph theoretical metrics, to discover the optimal entanglement generation path.
@@
It is worth noting that the employed cost metrics can be tailored to meet the requirements of specific experimental platforms, e.g., as we have done with the ‘emitter depth’ to address the limitations in emitters’ coherence time. Additionally, the application can be easily extended to the probabilistic fusion-based generation protocols, as in Ref. [5], by facilitating the expansion of both the size scale and diversity of resource states used as building blocks.